# Fair Resource Allocation in Multi-Task Learning

### v1

@@ 6 Applying $\alpha$-Fairness to Existing Methods
where $\alpha\in(-\infty,1)$ and $i\in[K]$. Here we omit the model parameter $\theta$ for simplicity. It can be observed that the gradient changes from $g_{i}$ to $g_{i}/l_{i}^{\alpha}$, where $\alpha$ controls the emphasis placed on tasks with different levels of difficulty. Take the example of simply summing $\alpha$-fair losses of all tasks
@@
If we choose $\alpha=0$, the objective is reduced to Linear Scalarization (LS) which minimizes the sum of all losses. If $\alpha\to 1$, the objective tends to minimize the sum of the logarithmic losses, which shares similarity with Scale-Invariant (SI). If $\alpha\to-\infty$, the objective exhibits the notion of the minimax fairness [33], which aims to minimize the maximum loss among all tasks.
Proposition 1.
The Pareto front of the $\alpha$-fair loss functions in eq. 5 is the same as that of original loss functions $(l_{1},...,l_{K})$.
According to 1, transforming each $l_{i}$ to its $\alpha$-fair counterpart does not change the Pareto front, and allows us to find an improved solution along this front under a proper selection of fairness.
We then apply this $\alpha$-fair loss transformation to a series of MTL methods including LS, RLW [61], DWA [37], UW [43], MGDA [23], PCGrad [20], CAGrad [21], and test the performance on NYU-v2 and Cityscapes datasets. We simply choose $\alpha=0.5$ for all the experiments of $\alpha$-fair losses. Other experiment settings remain the same with Section 5.1. The results presented in Table 6 and Table 7 clearly demonstrate that the $\alpha$-fair loss transformation improves the performance of these MTL methods via a large margin.
@@ 7 Theoretical Analysis
In this section, we provide a theoretical analysis of our method on the convergence to a Pareto stationary point, at which some convex combination of task gradients is 0. As mentioned before, we assume that the gradients of different tasks are linearly independent when not reaching a Pareto stationary point. Formally, we make the following assumption, as also adopted by [26].
Assumption 1.
@@
The following assumption imposes differentiability and Lipschitz continuity on the loss functions, as also adopted by [21; 26].
Assumption 2.
@@
Theorem 1.
Suppose Assumptions 1-2 are satisfied. Set the stepsize $\eta_{t}=\frac{\sum_{i}w_{t,i}^{-1/\alpha}}{LK\sum_{i}w_{t,i}^{1-1/\alpha}}$, Then, there exists a subsequence $\{\theta_{t_{j}}\}$ of the output sequence $\{\theta_{t}\}$ that converges to a Pareto stationary point $\theta^{\star}$.
@@
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU normal-$\uparrow$ Pix Acc normal-$\uparrow$ Abs Err normal-$\downarrow$ Rel Err normal-$\downarrow$ LS 75.18 93.49 0.0155 46.77 22.60 Fair-LS 74.91 93.48 0.0137 37.51 10.86 RLW 74.57 93.41 0.0158 47.79 24.38 Fair-RLW 74.32 93.36 0.0140 37.46 11.64 DWA 75.24 93.52 0.0160 44.37 21.45 Fair-DWA 75.06 93.46 0.0147 35.34 10.74 MGDA 68.84 91.54 0.0309 33.50 44.14 Fair-MGDA 74.45 93.50 0.0131 37.64 9.91 PCGrad 75.13 93.48 0.0154 42.07 18.29 Fair-PCGrad 75.25 93.51 0.0140 37.00 10.71 CAGrad 75.16 93.48 0.0141 37.60 11.64 Fair-CAGrad 74.74 93.39 0.0134 33.04 6.23
@@ 8 Conclusion
For the future study, we will explore the performance of FairGrad in more challenging MTL settings with significantly diverse tasks. Theoretically, we will study the impact of varying levels of difficulty across tasks on the final convergence and generalization performance.
Supplementary Materials
@@ A.1 Toy Example
Following [26, 28], we use a slightly modified version of the 2-task toy example provided in [21]. The two tasks $L_{1}(x)$ and $L_{2}(x)$ are defined on $x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}$,
@@ A.2 Detailed Results on Multi-Task Regression
Method $\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ MRnormal-$\downarrow$ $\Delta m\%\downarrow$ MAE normal-$\downarrow$ STL 0.067 0.181 60.57 53.91 0.502 4.53 58.8 64.2 63.8 66.2 0.072 LS 0.106 0.325 73.57 89.67 5.19 14.06 143.4 144.2 144.6 140.3 0.128 8.18 177.6 SI 0.309 0.345 149.8 135.7 1.00 4.50 55.3 55.75 55.82 55.27 0.112 4.82 77.8 RLW 0.113 0.340 76.95 92.76 5.86 15.46 156.3 157.1 157.6 153.0 0.137 9.55 203.8 DWA 0.107 0.325 74.06 90.61 5.09 13.99 142.3 143.0 143.4 139.3 0.125 7.82 175.3 UW 0.386 0.425 166.2 155.8 1.06 4.99 66.4 66.78 66.80 66.24 0.122 6.18 108.0 MGDA 0.217 0.368 126.8 104.6 3.22 5.69 88.37 89.4 89.32 88.01 0.120 7.73 120.5 PCGrad 0.106 0.293 75.85 88.33 3.94 9.15 116.36 116.8 117.2 114.5 0.110 6.36 125.7 CAGrad 0.118 0.321 83.51 94.81 3.21 6.93 113.99 114.3 114.5 112.3 0.116 7.18 112.8 IMTL-G 0.136 0.287 98.31 93.96 1.75 5.69 101.4 102.4 102.0 100.1 0.096 6.09 77.2 Nash-MTL 0.102 0.248 82.95 81.89 2.42 5.38 74.5 75.02 75.10 74.16 0.093 3.64 62.0 FAMO 0.15 0.30 94.0 95.2 1.63 4.95 70.82 71.2 71.2 70.3 0.10 4.73 58.5 FairGrad 0.117 0.253 87.57 84.00 2.15 5.07 70.89 71.17 71.21 70.88 0.095 3.82 57.9
@@ B.1 $\alpha$-fairness
Different values of $\alpha$ yield different ideas of fairness. Recall from eq. 1 the following utilization maximization objective
@@
When $\alpha\to 1$, the utilization maximization objective eq. 1 captures the proportional fairness. First note that
@@
For a concave function $f(x)$ over a domain $\mathcal{D}$, it is shown in [34] that
@@
Clearly, since the objective $\sum_{i\in[K]}\log x_{i}$ is concave, applying eq. 7 yields
@@
which corresponds to Nash-MTL [26].
When $\alpha\to\infty$, the utilization maximization objective eq. 1 yields the max-min fairness. The following proofs follow Section 2.2.1 in [34]. Let $x^{\star}(\alpha)$ be the $\alpha$-fair allocation. Assume $x_{i}^{\star}(\alpha)\to x^{\star}$ as $\alpha\to\infty$ and $x_{1}^{\star}<x_{2}^{\star}<\cdots<x_{K}^{\star}$. Let $\epsilon$ be the minimum difference of $\{x^{\star}\}$. That is, $\epsilon=\min_{i}|x_{i+1}^{\star}-x_{i}^{\star}|$, $i\in[K-1]$. When $\alpha$ is sufficiently large, we then have $|x_{i}^{\star}(\alpha)-x_{i}^{\star}|\leq\epsilon/4$, which also implies $x_{1}^{\star}(\alpha)<x_{2}^{\star}(\alpha)<\cdots<x_{K}^{\star}(\alpha)$. According to eq. 7, we have
@@ B.2 Convergence
Theorem 2 (Restatement of 1).
Suppose Assumptions 1-2 are satisfied. Set the stepsize $\eta_{t}=\frac{\sum_{i}w_{t,i}^{-1/\alpha}}{LK\sum_{i}w_{t,i}^{1-1/\alpha}}$, Then, there exists a subsequence $\{\theta_{t_{j}}\}$ of the output sequence $\{\theta_{t}\}$ that converges to a Pareto stationary point $\theta^{\star}$.
@@
For any fixed $K$, it can be concluded from eq. 8, eq. 9, and eq. 10 that
@@ B.3 $\alpha$-fair loss transformation
Proposition 2 (Restatement of 1).
The Pareto front of the $\alpha$-fair loss functions in eq. 5 is the same as that of original loss functions $(l_{1},...,l_{K})$.

### v2

@@ 6 Applying $\alpha$-Fairness to Existing Methods
where $\alpha\in(-\infty,1)$ and $i\in[K]$. Note that the meaning of $\alpha$ here differs from that used in Section 4. FairGrad aims to address the issue of varying loss decreasing rates, while applying $\alpha$-fairness to existing methods tries to deal with the issue of varying loss scales. Although the ideas of $\alpha$-fairness are the same, the goals are different.
Here we omit the model parameter $\theta$ for simplicity. It can be observed that the gradient changes from $g_{i}$ to $g_{i}/l_{i}^{\alpha}$, where $\alpha$ controls the emphasis placed on tasks with different levels of difficulty. Take the example of simply summing $\alpha$-fair losses of all tasks
@@
If we choose $\alpha=0$, the objective is reduced to Linear Scalarization (LS) which minimizes the sum of all losses. If $\alpha\to 1$, the objective tends to minimize the sum of the logarithmic losses, which shares similarity with Scale-Invariant (SI). If $\alpha\to-\infty$, the objective exhibits the notion of the minimax fairness (Radunovic & Le Boudec, 2007), which aims to minimize the maximum loss among all tasks.
Proposition 6.1.
The Pareto front of the $\alpha$-fair loss functions in Equation 5 is the same as that of original loss functions $(l_{1},...,l_{K})$.
According to Proposition 6.1, transforming each $l_{i}$ to its $\alpha$-fair counterpart does not change the Pareto front, and allows us to find an improved solution along this front under a proper selection of fairness.
We then apply this $\alpha$-fair loss transformation to a series of MTL methods including LS, RLW (Lin et al., 2021), DWA (Liu et al., 2019), UW (Kendall et al., 2018), MGDA (Sener & Koltun, 2018), PCGrad (Yu et al., 2020b), CAGrad (Liu et al., 2021), and test the performance on NYU-v2 and Cityscapes datasets. We simply choose $\alpha=0.5$ for all the experiments. Other experiment settings remain the same with Section 5.2. The results presented in Table 6 and Table 7 clearly demonstrate that the $\alpha$-fair loss transformation improves the performance of these MTL methods via a large margin.
Additionally, we also test the applicability of $\alpha$-fair loss transformation to FairGrad. It can be seen from Table 7 that compared to other MTL methods, applying this transformation to FairGrad provides only a marginal improvement. This shows that FairGrad can mitigate the issue of varying loss scales by incorporating fairness-based utility functions.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ LS 75.18 93.49 0.0155 46.77 22.60 Fair-LS 74.91 93.48 0.0137 37.51 10.86 RLW 74.57 93.41 0.0158 47.79 24.38 Fair-RLW 74.32 93.36 0.0140 37.46 11.64 DWA 75.24 93.52 0.0160 44.37 21.45 Fair-DWA 75.06 93.46 0.0147 35.34 10.74 MGDA 68.84 91.54 0.0309 33.50 44.14 Fair-MGDA 74.45 93.50 0.0131 37.64 9.91 PCGrad 75.13 93.48 0.0154 42.07 18.29 Fair-PCGrad 75.25 93.51 0.0140 37.00 10.71 CAGrad 75.16 93.48 0.0141 37.60 11.64 Fair-CAGrad 74.74 93.39 0.0134 33.04 6.23 FairGrad 75.72 93.68 0.0134 32.25 5.18 Fair-FairGrad 75.50 93.51 0.0131 32.54 4.92
@@ 7 Theoretical Analysis
In this section, we provide a theoretical analysis of our method on the convergence to a Pareto stationary point, at which some convex combination of task gradients is 0. As mentioned before, we assume that the gradients of different tasks are linearly independent when not reaching a Pareto stationary point. Formally, we make the following assumption, as also adopted by (Navon et al., 2022).
Assumption 7.1.
@@
The following assumption imposes differentiability and Lipschitz continuity on the loss functions, as also adopted by (Liu et al., 2021; Navon et al., 2022).
Assumption 7.2.
@@
Theorem 7.3.
Suppose Assumptions 7.1-7.2 are satisfied. Set the stepsize $\eta_{t}=\frac{\sum_{i}w_{t,i}^{-1/\alpha}}{LK\sum_{i}w_{t,i}^{1-1/\alpha}}$, Then, there exists a subsequence $\{\theta_{t_{j}}\}$ of the output sequence $\{\theta_{t}\}$ that converges to a Pareto stationary point $\theta^{\star}$.
@@ 8 Conclusion
For future studies, we will explore the performance of FairGrad in more challenging MTL settings with significantly diverse tasks. Theoretically, we will study the impact of varying levels of difficulty across tasks on the final convergence and generalization performance.
Impact Statement
This paper discusses the fairness in optimization methods for multi-task learning (MTL). There are some potential societal consequences, none of which we feel must be specifically highlighted here.
@@ A.1 Toy Example
Following (Navon et al., 2022; Liu et al., 2023), we use a slightly modified version of the 2-task toy example provided in (Liu et al., 2021). The two tasks $L_{1}(x)$ and $L_{2}(x)$ are defined on $x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}$,
@@ A.2 Detailed Results on Multi-Task Regression
[…]
A.3 Detailed Results on Effect of Different Fairness Criteria
We provide additional results of different fairness criteria discussed in Section 5.4 in Table 9 and Table 10. As $\alpha$ changes, the algorithm will prioritize certain tasks over others. Hence, an overall performance drop may be observed. However, tasks with lower priority may get higher ranks, then leading to a higher MR value.
Table 9: Results of Different Fairness Criteria on Cityscapes (2-task) dataset. Each experiment is repeated 3 times with different random seeds and the average is reported.
Method Segmentation Depth MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ FairGrad ($\alpha=1$) 75.94 93.65 0.0138 33.13 2.25 6.73 FairGrad ($\alpha=2$) 74.10 93.03 0.0135 29.92 5.75 3.90 FairGrad ($\alpha=5$) 67.30 90.23 0.0134 30.01 7.25 6.87 FairGrad ($\alpha=10$) 62.77 88.00 0.0151 28.05 8.50 10.54
Table 10: Results of Different Fairness Criteria on NYU-v2 (3-task) dataset. Each experiment is repeated 3 times with different random seeds and the average is reported.
Method Segmentation Depth Surface Normal MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 FairGrad ($\alpha=1$) 40.64 67.20 0.5671 0.2434 25.18 20.05 28.35 55.61 68.37 4.78 -2.79 FairGrad ($\alpha=2$) 38.80 65.29 0.5572 0.2322 24.55 18.97 30.50 57.94 70.14 4.78 -4.96 FairGrad ($\alpha=5$) 34.05 62.82 0.5853 0.2375 24.40 18.70 30.96 58.48 70.45 6.22 -3.03 FairGrad ($\alpha=10$) 31.45 61.43 0.5948 0.2341 24.80 19.08 30.10 57.58 69.69 6.22 -1.00
Table 11: Results of Different Methods for Solving Sub-problems of Nash-MTL on Cityscapes (2-task) dataset. Each experiment is repeated 3 times with different random seeds and the average is reported.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Nash-MTL (original) 75.41 93.66 0.0129 35.02 6.82 Nash-MTL (ours) 75.26 93.71 0.0129 34.45 6.28
A.4 Difference Between FairGrad and Nash-MTL
Although FairGrad with proportional fairness shares similarities with Nash-MTL, there are many differences. The high-level ideas are different. Nash-MTL is developed from the perspective of game theory, whereas our FairGrad is inspired by fair resource allocation in communication networks. Thus, this allows us to incorporate the advances from network resource allocation into MTL. FairGrad incorporates other different notions of fairness that Nash-MTL cannot cover. From our empirical studies on Cityscapes and NYUv2 datasets, the performance of MDP fairness is significantly better than proportional fairness. This indicates that proportional fairness may not always be the most suitable choice for different applications and scenarios. This greatly highlights the importance of incorporating other fairness ideas. Unlike Nash-MTL, our proposed FairGrad offers the flexibility to explore different fairness criteria.
Technically, algorithmic designs are different. We propose to solve weights ${w_{1},...,w_{K}}$ from our objective through a simple nonlinear least square problem. Solving this problem is efficient, and it turns out that the results are good enough. As a comparison, Nash-MTL solves weights from a different constrained objective via a variation of concave-convex-procedure (CCP) that solves a sequence of simple constrained problems. Our approach that solves a nonlinear least square equation can also be applied to Nash-MTL.
We experiment on the Cityscapes dataset using the codes from the Nash-MTL paper. The results are presented in Table 11, where Nash-MTL (original) denotes the original method, and Nash-MTL (ours) denotes the method using our approach to solve the sub-problems. The results demonstrated that our approach can not only be applied to Nash-MTL, but also achieve slightly better performance.
@@ B.1 $\alpha$-fairness
Different values of $\alpha$ yield different ideas of fairness. Recall from Equation 1 the following utilization maximization objective
$\max_{x_{1},...,x_{K}\in\mathcal{D}}\quad\sum_{i\in[K]}u(x_{i}):=\frac{x_{i}^{1-\alpha}}{1-\alpha},$
@@
When $\alpha\to 1$, the utilization maximization objective Equation 1 captures the proportional fairness. First note that
@@
For a concave function $f(x)$ over a domain $\mathcal{D}$, it is shown in (Srikant & Ying, 2013) that
@@
Clearly, since the objective $\sum_{i\in[K]}\log x_{i}$ is concave, applying Equation 6 yields
@@
which corresponds to Nash-MTL (Navon et al., 2022).
When $\alpha\to\infty$, the utilization maximization objective Equation 1 yields the max-min fairness. The following proofs follow Section 2.2.1 in (Srikant & Ying, 2013). Let $x^{\star}(\alpha)$ be the $\alpha$-fair allocation. Assume $x_{i}^{\star}(\alpha)\to x^{\star}$ as $\alpha\to\infty$ and $x_{1}^{\star}<x_{2}^{\star}<\cdots<x_{K}^{\star}$. Let $\epsilon$ be the minimum difference of $\{x^{\star}\}$. That is, $\epsilon=\min_{i}|x_{i+1}^{\star}-x_{i}^{\star}|$, $i\in[K-1]$. When $\alpha$ is sufficiently large, we then have $|x_{i}^{\star}(\alpha)-x_{i}^{\star}|\leq\epsilon/4$, which also implies $x_{1}^{\star}(\alpha)<x_{2}^{\star}(\alpha)<\cdots<x_{K}^{\star}(\alpha)$. According to Equation 6, we have
@@ B.2 Convergence
Theorem B.1 (Restatement of Theorem 7.3).
Suppose Assumptions 7.1-7.2 are satisfied. Set the stepsize $\eta_{t}=\frac{\sum_{i}w_{t,i}^{-1/\alpha}}{LK\sum_{i}w_{t,i}^{1-1/\alpha}}$, Then, there exists a subsequence $\{\theta_{t_{j}}\}$ of the output sequence $\{\theta_{t}\}$ that converges to a Pareto stationary point $\theta^{\star}$.
@@
For any fixed $K$, it can be concluded from Equation 7, Equation 8, and Equation 9 that
@@ B.3 $\alpha$-fair loss transformation
Proposition B.2 (Restatement of Proposition 6.1).
The Pareto front of the $\alpha$-fair loss functions in Equation 5 is the same as that of original loss functions $(l_{1},...,l_{K})$.